# Spread them Apart: Towards Robust Watermarking of Generated Content

### v1

@@ 4.2 Implementation Details
Here, $\varepsilon$ defines the minimum difference between private key pixels that we would like to obtain. Note that the larger the value of $\varepsilon$ is, the more robust the watermark is to additive perturbations. At the same time, the increase of $\varepsilon$ negatively influences the perceptual quality of images.
The second term $\mathcal{L}_{qual}$, is introduced to preserve the generation quality of the image. The value $\mathcal{L}_{qual}$ is a difference in image quality measured by LPIPS metric (Zhang et al. [2018]), which acts as a regularization. Given $x$ as the original image and $\hat{x}$ as the watermarked image, $\mathcal{L}_{qual}$ is defined as follows:
$\mathcal{L}_{qual}(x,\hat{x})=\sum_{j}\frac{1}{W_{j}H_{j}}\sum_{w,h}\|\phi^{j}(x)-\phi^{j}(\hat{x})\|^{2}_{2}.$ (12)
Here, $\phi^{j}(x)=w_{j}\odot o^{j}_{hw}(x)$, where $o^{j}(x)$ are the internal activations of the CNN, AlexNet (Krizhevsky et al. [2012]), in our case.
Note that we do not perform denoising at each iteration, as we only manipulate the latent vectors produced by U-Net; the forward step of the described optimization procedure involves only the decoding of the latent vectors: $x=\mathcal{D}(z)$.
The optimization is performed over $700$ steps of the Adam optimizer with the learning rate of $8\times 10^{-3}$, where every $100$ iteration, the learning rate is halved. When the convergence is reached, the ordinary Stable Diffusion post-processing of the image is performed. The coefficients $\lambda_{wm}$ and $\lambda_{qual}$ are determined experimentally and set to be $0.9$ and $150$, respectively, the value of $\varepsilon$ was set to be $\varepsilon=0.2$. Schematically, the process of watermark embedding and extraction is presented in Figure 1.
4.3 Robustness Guarantee
By construction, the watermark embedded by our method is robust against additive watermark removal attacks of a bounded magnitude. Namely, let the watermark $w(u_{i}|x)$ be embedded in $x$ with the use of the secret $s(u_{i})=\{a_{1},\dots,a_{n},b_{1},\dots,b_{n}\}$ of the user $u_{i}$. Let
$\Delta_{i}=\frac{|x_{a_{i}}-x_{b_{i}}|}{2}.$ (13)
Then, the following lemma holds.
Lemma 4.1.
Let $\varepsilon\in\mathbb{R}^{d}$ and $\Delta_{i_{1}}\leq\Delta_{i_{2}}\leq\dots\leq\Delta_{i_{n}}$.
Then, if $\|\varepsilon\|_{\infty}<\Delta_{i_{k}}$, then $d(w(u_{i}|x+\varepsilon),w(u_{i}|x))<k.$
Proof.
Note that to change the $j^{\prime}$th bit of watermark $w(u_{i}|x)$, an adversary has to change the sign in expression $(x_{a_{j}}-x_{b_{j}})$. Without the loss of generality, let $x_{a_{j}}-x_{b_{j}}\geq 0$.
Consider an additive noise $\varepsilon$ such that $(x+\varepsilon)_{a_{j}}-(x+\varepsilon)_{b_{j}}<0,$ meaning $|\varepsilon_{b_{j}}-\varepsilon_{a_{j}}|>|x_{a_{j}}-x_{b_{j}}|$. Note that $\|\varepsilon\|_{\infty}\geq\max(|\varepsilon_{a_{j}}|,|\varepsilon_{b_{j}}|).$
If $\max(|\varepsilon_{a_{j}}|,|\varepsilon_{b_{j}}|)<\Delta_{j},$ then $|\varepsilon_{b_{j}}-\varepsilon_{a_{j}}|\leq|\varepsilon_{b_{j}}|+|\varepsilon_{a_{j}}|<2\Delta_{j}=|x_{a_{j}}-x_{b_{j}}|,$ yielding a contradiction. Thus, $\|\varepsilon\|_{\infty}\geq\Delta_{j}.$
Finally, an observation that all the indices in $s(u_{i})$ are unique finalizes the proof. ∎
This lemma provides a lower bound on the $l_{\infty}$ norm of the additive perturbation $\varepsilon$ applied to $x$ which is able to erase at least $k$ bits of the watermark $w(u_{i}|x))$ embedded into $x$.
@@ 5.1 General Setup
We use stable-diffusion-2-base model (Rombach et al. [2022]) with the epsilon prediction type and $50$ steps of denoising for the experiments. The resolution of generated images is $512\times 512$. The experiments were conducted on DiffusionDB dataset (Wang et al. [2022]). Specifically, we choose $1000$ unique prompts and generate $1000$ different images.
The public key for the user is sampled from the Bernoulli distribution with the parameter $p=0.5$. The length of a key is set to be $n=100$. The private key is generated by randomly picking $2n$ unique pairs of indices of the flattened image.
@@ 5.2 Attack Details
We evaluate the robustness of the watermarks embedded by our method against the following watermark removal attacks: brightness adjustment, contrast shift, gamma correction, image sharpening, hue adjustment, saturation adjustment, random additive noise, JPEG compression, and the white-box PGD adversarial attack (Madry et al. [2018]). In this section, we describe these attacks in detail.
@@
On the contrary, when the contrast shift is performed with the negative value of $c$ (namely, $c\sim\mathcal{U}[-2,-0.5]$), such a transform turns an image into a negative. Later, we treat these transforms separately and denote them as “Contrast $+$” and “Contrast --”, depending on the sign of $c$.
@@
For sharpening, hue, and saturation adjustment, we use implementations from the Kornia package (Riba et al. [2020]) with the following parameters: $a_{saturation}=2,\quad a_{hue}=0.2,$ $\quad a_{sharpness}=2.$
The noise for the noising attack was sampled from the uniform distribution $\mathcal{U}[-\delta,\delta]$, where $\delta=25$. Note, that the maximum $\|\cdot\|_{\infty}$ of noise is then equal to $25$.
JPEG compression was performed using DiffJPEG (Shin [2017]) with quality equal to $50$.
White-box attack aims to change the embedded watermark $w$ to some other watermark $\tilde{w}$ by optimizing the image with respect to the loss initially used to embed the watermark $w$:
@@
$\mathcal{L}_{wm}=\sum_{i=1}^{n}\min((-1)^{\tilde{w}_{i}}(x_{a_{i}}-x_{b_{i}})+\varepsilon,0).$
In Eq. 14, the term $\mathcal{L}_{qual}$ corresponds to the difference in image quality in terms of LPIPS metric, namely,
$\mathcal{L}_{qual}=LPIPS(x,\hat{x}),$ (15)
where $x$ and $\hat{x}$ are the original image and image on a particular optimization iteration, respectively.
The loss function $\mathcal{L}_{wb}$ pushes the private key pixels to be aligned with a new randomly sampled public key $\tilde{w}$ so that the ground-truth watermark $w$ gets erased. The attack’s budget is the upper bound of $\|\cdot\|_{\infty}$ norm of the additive perturbation, that we have taken to be ${\varepsilon}/{2}$ from the Eq. 11. Let $\tilde{x}$ be the image obtained after the attack. If at some iteration the distance between the source image $x$ and the attacked one $\tilde{x}$ exceeds ${\varepsilon}/{2}$, $\tilde{x}$ is being projected back onto the sphere $\|\tilde{x}-x\|_{\infty}={\varepsilon}/{2}$. The optimization took place for $10$ iterations with the Adam optimizer and the learning rate was equal to $10^{-1}$. Note that this attack setting implies knowledge about the private key and assumes white-box access to the generative model. Hence, this is de facto the strongest watermark removal attack we consider.
Pixels of the images perturbed by the attacks are then linearly mapped to $[0,255]$ segment:
$x^{(i)}=255\frac{x^{(i)}-x^{(i)}_{min}}{x^{(i)}_{max}-x^{(i)}_{min}},\quad i\in\{R,G,B\}.$ (16)
5.3 Results
In this section, we provide the quantitative results of experiments. We report (i) quality metrics of the generated images (SSIM, PSNR, FID and LPIPS) to evaluate the invisibility of the watermarks, (ii) bit-wise error of the watermark extraction caused by watermark removal attacks and (iii) True Positive Rates in attribution and detection problems.
We compare our results (where applicable) to that of Stable Signature Fernandez et al. [2023], SSL watermarking Fernandez et al. [2022], AquaLora Feng et al. [2024] and WOUAF Kim et al. [2024], one of the state-of-the-art watermarking approaches. In these works the watermark length is set to be $48$, $30$, $48$, and $32$, respectively, while we have $100$ bits long watermarks: note that the longer the watermark, the harder it is to be embedded.
The image quality metrics are presented in Table 4. It can be seen that our results are comparable to the ones of the baseline methods in terms of the quality of the produced images. A qualitative comparison of original and watermarked images can be found in Figure 2. To evaluate the robustness of the watermarks against removal attacks, we report an average bit-wise error, ABWE:
@@ 5.3 Results
where $w^{gt}_{i,j}$ and $w^{extracted}_{i,j}$ are the $j$-th bits of ground truth and extracted private keys, corresponding to the $i$-th image. Here, $n$ is the number of bits in the watermark and $N$ is the number of images. We report ABWE in Table 1.
Table 1: Average bit-wise error after watermark removal attacks. The column “Generation” corresponds to the average bit-wise error of the watermarking process. The best results are highlighted in bold.
Method Generation Brightness Contrast $+$ Contrast -- Gamma JPEG
Ours $0.0008$ 0.002 0.002 0.998 0.003 $0.147$
Stable signature $0.01$ $0.03$ $0.02$ $0.12$
SSL watermarking 0.00 $0.06$ $0.04$ 0.04
AquaLora $0.0721$ $0.137$ $0.137$ $0.0508$
@@
Ours 0.01 $0.1$ 0.0008 0.057 0.064
Stable signature 0.01 $0.01$
SSL watermarking $0.06$
AquaLora $0.137$ $0.137$ $0.07$
Table 2: TPRs under different types of watermark removal attacks, attribution problem. We use $k=10$ different private keys and fix FPR = $10^{-6}$. Such a FPR is achieved when $\tau_{1}=25$ and $\tau_{2}=75$ from Eq. 5. The parameters of removal attacks are presented in Section 5.2. The best results are highlighted in bold.
Method Generation Brightness Contrast $+$ Contrast -- Gamma JPEG
Ours 1.000 1.000 1.000 1.000 1.000 $0.444$
Stable signature $0.998$ $0.927$ $0.984$ $0.784$
AquaLora $0.998$ $0.941$ $0.941$ 0.998
WOUAF 1.000 $0.975$ $0.969$
@@
Ours 1.000 0.6530.653{0.653}0.653 1.000 0.971 0.862
Stable signature 0.998 $0.776$ $0.747$
AquaLora $0.941$ 0.9410.941{0.941}0.941 $0.958$
WOUAF 0.9700.970{0.970}0.970
Table 3: TPRs under different types of watermark removal attacks, detection problem. We use $k=10$ different private keys and fix FPR = $10^{-6}$. Such a FPR is achieved when $\tau_{1}=25$ and $\tau_{2}=75$ from Eq. 5. The parameters of removal attacks are presented in Section 5.2. The best results are highlighted in bold.
Method Generation Brightness Contrast $+$ Contrast -- Gamma JPEG
Ours 1.000 1.000 1.000 1.000 1.000 $0.444$
Stable signature 1.000 $0.862$ $0.984$ $0.217$
SSL watermarking $1.000$ $0.940$ $0.960$ 0.810
@@
Ours 1.000 0.6530.653{0.653}0.653 1.000 0.971 0.862
Stable signature 0.998 $0.406$ $0.505$
SSL watermarking 1.000
To estimate the TPR in the attribution problem, we extract $k=10$ different watermarks from the watermarked images. To extract a different watermark, we randomly generate $k=10$ different private keys to simulate other users. The results are reported in Table 2 together with the TPRs under different watermark removal attacks. Note that the PGD attack in this setting is aimed at restoring the original watermark. To estimate the TRP in the watermark detection problem, we do the same procedure for non-watermarked images generated by the Stable Diffusion model and extract $k=10$ different watermarks. The results are presented in Table 3.
Note that our framework yields both low misattribution and misdetection rates according to the two-tail detection and attribution rules from the Eq. 9. The proposed approach yields watermarks that are provably robust to additive perturbations of a bounded magnitude, multiplicative perturbations of any kind, and exponentiation.
5.4 Increasing the Robustness to Watermark Removal Attacks
It is noteworthy that the set of watermark removal attacks described in Section 5.2 does not include geometric transformations. The version of the proposed method described in Section 4 does not provide robustness against geometric transformations of the image, however, the method can be extended to support them. In Section A of the supplementary, we discuss an extension of the proposed method to provide robustness against rotations, and translations. An extension is based on a simultaneous embedding of the watermark in the pixel space and special functions, which are invariant to certain transformations. An intuition behind this extension is as follows. Suppose that the transform $\gamma:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is invariant under parametric perturbation $\phi:\mathbb{R}^{d}\times\Theta\to\mathbb{R}^{d}$, namely, $\gamma(x)=\gamma(\phi(x,\theta))$ for all $x\in\mathbb{R}^{d},\ \theta\in\Theta$. Then, if the watermark is successfully embedded into $\gamma(x)$ it becomes robust under perturbation $\phi$.
Table 4: Image quality metrics. The best results are highlighted in bold.
Metric St. Sign. AquaLora WOUAF Ours
[…]
PSNR $\uparrow$ 30.0 $29.42$ $29.4$
FID $\downarrow$ $19.6$ $24.72$ $>15.0$ 13.2
LPIPS $\downarrow$ $0.0072$
Original image Watermarked image Pixel-wise difference (×$\times 10$)
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 2: Examples of watermarked images. The maps of absolute pixel-wise difference between source images and the generated ones were added for illustrational purposes.
6 Conclusion and Future Work
In this paper, we propose Spread them Apart, the framework to watermark generated content of continuous nature and apply it to images generated by Stable Diffusion. We prove that the watermarks produced by our method are provably robust against additive watermark removal attacks of a bounded norm and are provably robust to multiplicative perturbations by design. Our approach can be used to both detect that the image is generated by a given model and to identify the end-user who generated it. Experimentally, we show that our method is comparable to the state-of-the-art watermarking methods in terms of the invisibility of watermark and the robustness to synthetic watermark removal attacks.
Appendix A Improving the Robustness to Watermark Removal Attacks
Note that the version of the watermarking method proposed in this paper is based on the embedding of the watermark into the pixel domain of a generated image. Hence, the method can not yield watermarks provably robust against geometric transformations of the image, such as rotations and translations. In this section, we describe an extension of the method to guarantee robustness against such transformations. An extension is done by embedding the watermark simultaneously into the pixel domain and special representations in the frequency domain, which are invariant under rotations and translations.
A.1 Invariants in the frequency domain
where $\Theta$ is the set of parameters of perturbation $\phi$. In this work, we use two invariants discussed in Lin and Brandt [1993], formulated as theorems below.
Theorem A.1.
Let $h(x,y)$ be an integrable nonnegative function and its Fourier transform
[…]
be twice differentiable. Then the function $A(\omega_{x},\omega_{y})$ is invariant under translation.
Theorem A.2.
[…]
where $\tilde{A}(\omega,k)$ is the magnitude and $\tilde{\psi}(\omega,k)$ is the phase.
If $\tilde{h}(r,t)$ is an integrable nonnegative function and its Fourier-Mellin transform $\tilde{H}(\omega,k)$ is twice differentiable, then the function $\tilde{A}(\omega,k)$ is invariant under rotation.
Remark.
For consistent notation, we assume that the image $h(x,y)$ is a scalar function on a two-dimensional plane. Later, we refer to invariants from Theorems A.1-A.2 as to $\gamma_{t}$ and $\gamma_{r}$, respectively.
A.2 Spread Them Apart: Three Watermarks Instead of One
$w(u_{i})=\{w_{1},\dots,w_{n}\}$ is the watermark assigned to user $u_{i}$ and $s(u_{i})=\{a_{1},\dots,a_{n},b_{1},\dots,b_{n}\}$ is the secret of user $u_{i}$.
To ensure the robustness of the watermark to geometric transformations, we suggest embedding the watermark simultaneously in the pixel domain and in invariants $\gamma_{t}$ and $\gamma_{r}$. To do so, we optimize the loss function $\tilde{\mathcal{L}}$ in the form below:
~$\tilde{\mathcal{L}}=\lambda_{wm}\mathcal{L}_{wm}+\lambda_{qual}\mathcal{L}_{qual}+\lambda_{t}\mathcal{L}_{t}+\lambda_{r}\mathcal{L}_{r}=\mathcal{L}+\lambda_{t}\mathcal{L}_{t}+\lambda_{r}\mathcal{L}_{r}.$ (22)
In Eq. 22, $\lambda_{t},\ \lambda_{r}$ are positive constants and
$\mathcal{L}_{t}=\sum_{i=1}^{n}\min((-1)^{w_{i}}(\gamma_{t}(x)_{a_{i}}-\gamma_{t}(x)_{b_{i}})+\varepsilon,0),$ (23)
$\mathcal{L}_{r}=\sum_{i=1}^{n}\min((-1)^{w_{i}}(\gamma_{r}(x)_{a_{i}}-\gamma_{r}(x)_{b_{i}})+\varepsilon,0)$ (24)
are the loss functions that control the embedding of the watermark into invariants $\gamma_{t}(x)$ and $\gamma_{r}(x)$, respectively.
A.3 Extractions of watermarks
$\begin{cases}\gamma_{r}(x)_{a_{j}}\geq\gamma_{r}(x)_{b_{j}}\Longrightarrow w(u_{i}|\gamma_{r}(x))_{j}=0,\\ \gamma_{r}(x)_{a_{j}}<\gamma_{r}(x)_{b_{j}}\Longrightarrow w(u_{i}|\gamma_{r}(x))_{j}=1,\end{cases}$ (26)
$\begin{cases}\gamma_{t}(x)_{a_{j}}\geq\gamma_{t}(x)_{b_{j}}\Longrightarrow w(u_{i}|\gamma_{t}(x))_{j}=0,\\ \gamma_{t}(x)_{a_{j}}<\gamma_{t}(x)_{b_{j}}\Longrightarrow w(u_{i}|\gamma_{t}(x))_{j}=1.\end{cases}$ (27)
To assign the (possibly) watermarked image $x$ to the user, the owner of the model determines three candidates:
$\begin{cases}u=\arg\min\limits_{u_{i}\in[u_{1},\dots,u_{m}]:\xi_{i}=1}d(w(u_{i}),w(u_{i}|x)),\\ u_{r}=\arg\min\limits_{u_{i}\in[u_{1},\dots,u_{m}]:\xi^{r}_{i}=1}d(w(u_{i}),w(u_{i}|\gamma_{r}(x))),\\ u_{t}=\arg\min\limits_{u_{i}\in[u_{1},\dots,u_{m}]:\xi^{t}_{i}=1}d(w(u_{i}),w(u_{i}|\gamma_{t}(x))),\end{cases}$ (28)
where
indicates which of the users’ watermarks are within an appropriate distance from the corresponding extracted watermarks.
$\tilde{u}=\arg\min\{d(w(u),w(u|x)),\ d(w(u_{r}),w(u_{r}|\gamma_{r}(x))),\ d(w(u_{t}),w(u_{t}|\gamma_{t}(x)))\}.$ (30)
A.4 Probability of Incorrect Attribution
Assume that the user $u_{i}$ owns the watermarked image $x$. Note that the attribution of the image to the user $u_{i}$ is guaranteed to hold if
$u=u_{i},u_{r}=u_{i},u_{t}=u_{i}.$ (31)
Hence, the probability of incorrect attribution, ${FPR}_{3}(m)$, is bounded from above by the sum
$\mathbb{P}(u\neq u_{i})+\mathbb{P}(u_{r}\neq u_{i})+\mathbb{P}(u_{t}\neq u_{i}),$ (32)
yielding
$FRP_{3}(m)\leq 3\hat{p},$ (33)
where $\hat{p}$ is from Eq. LABEL:eq:fpr_m.
A.5 Quantitative Results
In Tables 5-6, we include rotation and translation transformations of the image. To rotate an image, we sample an angle of rotation $\theta_{r}$ uniformly from $[-10^{\circ},10^{\circ}]$; to translate an image, we sample the vector of translation $(\theta^{x}_{t},\theta^{y}_{t})$ uniformly from the set $[-10^{\circ},10^{\circ}]\times[-10^{\circ},10^{\circ}]$. We compare the results of an extended version of the proposed method with the baseline approach. STA(1) refers to the baseline approach described in Section 4, STA(3) refers to an extended approach described in Section A. Note that now, according to Section A.4, to achieve FPR $=10^{-6}$ we set $\tau_{1}=24$ and $\tau_{2}=76$. In the column “Error”, we report the bit-wise accuracy of the watermarking procedure.
To focus on the evaluation of the robustness of the watermarks, we report results for the images in which the watermark is embedded successfully. Namely, after embedding the watermark, we choose the watermarked images with a generation error smaller than $0.05$.
It is noteworthy that the simultaneous embedding of the watermark into the pixel domain and corresponding invariants in the frequency domain significantly improves the robustness against post-processing watermark removal attacks. The limitation of this approach is the increased complexity of the watermark embedding process which is the subject of future work.
Table 5: TPRs under different types of watermark removal attacks, attribution problem. We use $k=10$ different private keys and fix FPR = $10^{-6}$. The best results are highlighted in bold.
Method Error Brightness Contrast $+$ Contrast -- Gamma JPEG Translation
STA(1) 0.9720.972{0.972}0.972 0.9280.928{0.928}0.928 0.9280.928{0.928}0.928 0.9280.928{0.928}0.928 0.9280.928{0.928}0.928 0.724 $0.000$
STA(3) 0.980 1.000 1.000 1.000 0.951 0.5630.563{0.563}0.563 0.962
@@ A.5 Quantitative Results
Table 6: TPRs under different types of watermark removal attacks, detection problem. We use $k=10$ different private keys and fix FPR = $10^{-6}$. The best results are highlighted in bold.
Method Error Brightness Contrast $+$ Contrast -- Gamma JPEG Translation
STA(1) 0.9720.972{0.972}0.972 0.9280.928{0.928}0.928 0.9280.928{0.928}0.928 0.9280.928{0.928}0.928 0.9280.928{0.928}0.928 0.724 $0.000$
STA(3) 0.980 1.000 1.000 1.000 0.951 0.5630.563{0.563}0.563 0.962

### v2

@@ 5.2 Three Watermarks Instead of One
$w(u_{i})=\{w_{1},\dots,w_{n}\}$ is the watermark assigned to user $u_{i}$ and $s(u_{i})$ is the secret of user $u_{i}$. To ensure the robustness of the watermark to geometric transformations, we suggest embedding the watermark simultaneously in the pixel domain and in invariants $\gamma_{t}$ and $\gamma_{r}$. To do so, we optimize the loss function $\tilde{\mathcal{L}}$ in the form below:
$\tilde{\mathcal{L}}=\lambda_{wm}\mathcal{L}_{wm}+\lambda_{qual}\mathcal{L}_{qual}+\lambda_{t}\mathcal{L}_{t}+\lambda_{r}\mathcal{L}_{r}=\mathcal{L}+\lambda_{t}\mathcal{L}_{t}+\lambda_{r}\mathcal{L}_{r}.$ (16)
In Eq. 16, $\lambda_{t},\ \lambda_{r}$ are positive constants and $\mathcal{L}_{p}=\sum_{i=1}^{n}\min((-1)^{w_{i}}(\gamma_{p}(x)_{a_{i}}-\gamma_{p}(x)_{b_{i}})+\varepsilon,0)$ is the loss function that controls the embedding of the watermark into invariants, $\gamma_{t}(x)$ for $p=t$ and $\gamma_{r}(x)$ for $p=r$, respectively.
5.3 Extractions of Watermarks
$\begin{cases}\gamma_{p}(x)_{a_{j}}\geq\gamma_{p}(x)_{b_{j}}\Longrightarrow w(u_{i}|\gamma_{p}(x))_{j}=0,\\ \gamma_{p}(x)_{a_{j}}<\gamma_{p}(x)_{b_{j}}\Longrightarrow w(u_{i}|\gamma_{p}(x))_{j}=1,\end{cases}$ (18)
for $p=t$ and $p=r$ simultaneously. To assign the (possibly) watermarked image $x$ to the user, the owner of the model determines three candidates:
$\begin{cases}u=\arg\min\limits_{u_{i}\in[u_{1},\dots,u_{m}]:\xi_{i}=1}d(w(u_{i}),w(u_{i}|x)),\\ u_{p}=\arg\min\limits_{u_{i}\in[u_{1},\dots,u_{m}]:\xi^{p}_{i}=1}d(w(u_{i}),w(u_{i}|\gamma_{p}(x))),\end{cases}$ (19)
where the variables
indicate which of the users’ watermarks are within an appropriate distance from the corresponding extracted watermarks, given the invariant $p$.
$\tilde{u}=\arg\min$ $\{d(w(u),w(u|x)),d(w(u_{r}),w(u_{r}|\gamma_{r}(x))),d(w(u_{t}),w(u_{t}|\gamma_{t}(x)))\}.$
5.4 Probability of Incorrect Attribution
Assume that the user $u_{i}$ owns the watermarked image $x$. Note that the attribution of the image to the user $u_{i}$ is guaranteed to hold if $u=u_{i},u_{r}=u_{i},u_{t}=u_{i}.$ Hence, the probability of incorrect attribution, ${FPR}_{3}(m)$, is bounded from above by the sum $\mathbb{P}(u\neq u_{i})+\mathbb{P}(u_{r}\neq u_{i})+\mathbb{P}(u_{t}\neq u_{i}),$ yielding $FRP_{3}(m)\leq 3\hat{p}$, where $\hat{p}$ is from Eq. 4.
@@ 6.1 General Setup
We use stable-diffusion-2-base model [rombach2022high] with the epsilon prediction type and $50$ steps of denoising for the experiments. The resolution of generated images is $512\times 512$. The experiments were conducted on DiffusionDB dataset [wangDiffusionDBLargescalePrompt2022]. Specifically, we choose $1000$ unique prompts and generate $1000$ different images. The public key for the user is sampled from the Bernoulli distribution with the parameter $p=0.5$. The length of a key is set to be $n=100$. The private key is generated by randomly picking $n$ unique pairs of indices of the flattened image.
@@ 6.2 Attack Details
We evaluate the robustness of the watermarks embedded by our method against the following watermark removal attacks: brightness adjustment, contrast shift, gamma correction, image sharpening, hue adjustment, saturation adjustment, random additive noise, JPEG compression, and the white-box PGD adversarial attack [madry2018towards]. In this section, we describe these attacks in detail.
@@
  • On the contrary, when the contrast shift is performed with the negative value of $c$ (namely, $c\sim\mathcal{U}[-2,-0.5]$), such a transform turns an image into a negative. Later, we treat these transforms separately and denote them as “Contrast ++” and “Contrast -”, depending on the sign of $c$. Note that “Contrast -” imitates a bit flipping attack.
@@
  • For sharpening, hue, and saturation adjustment, we use implementations from the Kornia package [riba2020kornia] with the following parameters: $a_{saturation}=2,\quad a_{hue}=0.2,$ $\quad a_{sharpness}=2.$
  • The additive noise was sampled from the uniform distribution $\mathcal{U}[-\delta,\delta]$, where $\delta=25$. Note that the maximum $\|\cdot\|_{\infty}$ of noise is then equal to $25$.
  • JPEG compression was performed using DiffJPEG [Shin2017JPEGresistantAI] with quality equal to $50$.
  • White-box attack aims to change the embedded watermark $w$ to some other watermark $\hat{w}$ by optimizing the image with respect to the loss initially used to embed the watermark $w$:
@@
    $\mathcal{L}_{wm}=\sum_{i=1}^{n}\min((-1)^{\hat{w}_{i}}(x_{a_{i}}-x_{b_{i}})+\varepsilon,0).$
[…]
    The loss function $\mathcal{L}_{wb}$ pushes the private key pixels to be aligned with a new randomly sampled public key $\tilde{w}$ so that the ground-truth watermark $w$ gets erased. The attack’s budget is the upper bound of $\|\cdot\|_{\infty}$ norm of the additive perturbation, which we have taken to be ${\varepsilon}/{2}$ (see Eq. 11).
    If at some iteration the distance between the source image $x$ and the attacked one $\hat{x}$ exceeds ${\varepsilon}/{2}$, then we project $\hat{x}$ back onto the sphere $\|\hat{x}-x\|_{\infty}={\varepsilon}/{2}$. The optimization is performed for $10$ iterations with the Adam optimizer and the learning rate of $10^{-1}$. Note that this attack setting implies knowledge about the private key and assumes white-box access to the generative model. Hence, this is de facto the strongest watermark removal attack we consider.
Pixels of the images perturbed by the attacks are then linearly mapped to $[0,255]$ segment.
6.3 Results: Pixel Level Watermarking
We report (i) bit-wise error of the watermark extraction caused by watermark removal attacks and (ii) True Positive Rates in attribution and detection problems. We compare our results to those of Stable Signature [fernandez2023stable], SSL watermarking [fernandez2022watermarking], AquaLora [feng2024aqualora], and WOUAF [kim2024wouaf], one of the state-of-the-art image watermarking approaches. In these works, the watermark length is set to be $48$, $30$, $48$, and $32$, respectively, while we have $100$ bits long watermarks; note that our method allows embedding of the longest watermarks, among the considered competitors. A qualitative comparison of original and watermarked images can be found in the supplementary material. To evaluate the robustness of the watermarks against removal attacks, we report an average bit-wise error, ABWE:
@@ 6.3 Results: Pixel Level Watermarking
where $w^{gt}_{i,j}$ and $w^{extracted}_{i,j}$ are the $j$-th bits of ground truth and extracted watermarks, corresponding to the $i$-th image. Here, $n$ is the watermark length and $N=1000$ is the number of images. We report ABWE in Table 1.
Table 1: Average bit-wise error after watermark removal attacks. The column “Error” corresponds to the average bit-wise error of the watermarking process.
Method Error Brightness Contrast ++ Contrast - Gamma JPEG
Ours $0.001$ 0.002 0.002 0.002 0.003 $0.147$
Stable signature $0.012$ $0.025$ $0.025$ $0.518$ $0.016$ $0.167$
SSL 0.000 $0.117$ $0.071$ $0.785$ $0.041$ 0.205{0.205}
AquaLora $0.043$ $0.043$ $0.043$ $0.544$ $0.049$ 0.056{0.056}
WOUAF $0.031$ $0.005$ $0.008$ $0.317$ $0.004$ 0.014
@@
Ours 0.010 $0.110$ 0.001 0.056{0.056} 0.064{0.064}
Stable signature $0.014$ 0.014 $0.010$ $0.136$ $0.487$
SSL $0.061$ $0.198$ $0.003$ $0.254$ $0.272$
AquaLora $0.047$ $0.059$ $0.049$ $0.073$ 0.052
WOUAF $0.010$ $0.041$ $0.005$ 0.005 $0.132$
Table 2: TPRs under different watermark removal attacks, attribution problem.
Method Hit Brightness Contrast ++ Contrast - Gamma JPEG
Ours 1.000 1.000 1.000 1.000 1.000 $0.452$
Stable signature $0.983$ $0.977$ $0.976$ $0.000$ $0.982$ $0.576$
SSL $1.000$ $0.859$ $0.905$ $0.666$ $0.996$ $0.673$
AquaLora $0.955$ $0.955$ $0.953$ $0.178$ $0.940$ 0.936{0.936}
WOUAF 0.999{0.999} $0.997$ $0.993$ $0.002$ $0.995$ 0.967
@@
Ours 1.000 0.653{0.653} 1.000 0.971{0.971} 0.993
Stable signature $0.982$ 0.982 $0.983$ $0.675$ $0.000$
SSL $0.987$ $0.625$ $0.996$ $0.490$ $0.449$
AquaLora $0.949$ 0.914{0.914} $0.941$ $0.908$ $0.947$
WOUAF $0.999$ $0.862$ $0.997$ 0.974 $0.197$
Table 3: TPRs under different watermark removal attacks, detection problem.
Method Hit Brightness Contrast ++ Contrast - Gamma JPEG
Ours 1.000 1.000 1.000 1.000 1.000 $0.452$
Stable signature 0.983{0.983} $0.977$ $0.976$ $0.000$ $0.982$ $0.576$
SSL $1.000$ $0.863$ $0.906$ $0.674$ $0.996$ 0.681{0.681}
AquaLora $0.955$ $0.955$ $0.953$ $0.178$ $0.940$ 0.936{0.936}
WOUAF 0.999{0.999} $0.997$ $0.993$ $0.002$ $0.995$ 0.967
@@
Ours 1.000 0.653{0.653} 1.000 0.971{0.971} 0.993
Stable signature $0.982$ 0.982 $0.993$ $0.675$ $0.000$
SSL 0.987{0.987} $0.628$ $0.996$ $0.495$ $0.459$
AquaLora $0.949$ 0.914{0.914} $0.941$ $0.908$ $0.947$
WOUAF $0.999$ $0.862$ $0.997$ 0.974 $0.197$
Table 4: TPRs under different watermark removal attacks, attribution problem.
Method Hit Brightness Contrast ++ Contrast - Gamma JPEG Translation
STA(1) 0.972{0.972} 0.928{0.928} 0.928{0.928} 0.928{0.928} 0.928{0.928} 0.724 $0.000$
STA(3) 0.980 1.000 1.000 1.000 0.951 0.563{0.563} 0.962
@@
Table 5: TPRs under different watermark removal attacks, detection problem.
Method Hit Brightness Contrast ++ Contrast - Gamma JPEG Translation
STA(1) 0.972{0.972} 0.928{0.928} 0.928{0.928} 0.928{0.928} 0.928{0.928} 0.724 $0.000$
STA(3) 0.980 1.000 1.000 1.000 0.951 0.563{0.563} 0.962